Solve the equation and check your solution.8x + 3(x - 2) = 11x + 7
Answer by lwsshak3(11628)

Solve the equation and check your solution.
8x + 3(x - 2) = 11x + 7
8x+3x-6=11x+7
11x-11x=13
0≠13
no solution